The lenght of a rectangle is 5 inches more than the width. The perimeter is 34 inches. Find the lenght and the width?
Answer by mananth(12270)

width be x
length = x+5
..
Perimeter P = 2*(L+W)
2(x+x+5)=34
4x+10=34
divide by 2
2x+5=17
-5
2x= 17-5
2x=12
x=6in
length = x+5= 6+5 =11 in